OpenStudy (anonymous):

MODELING - Complementary Angles The difference between the measures of two complementary angles is 16°. Determine the measures of the two angles.

OpenStudy (anonymous):

complementary angles are angles whose sum is 90 let x, y be two complementary angles x+y = 90 -------- (1) x-y = 16 -------- (2)

OpenStudy (anonymous):

you can solve them i hope ?

OpenStudy (anonymous):

x+y=90 x-y=16 add them: 2x = 106 x = 53 Get y: y = 90-53 = 37 They are 37 degrees and 53 degrees